# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ 3.1 Generalization error analysis for calibration
We show the first PAC-Bayes bound for the generalization error of the ECE as follows.
@@
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. For both binary and multiclass settings, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (7)
Proof sketch.
We describe the case for binary classification. We first derive the following reformulation: $\mathrm{ECE}(f_{w},S_{\mathrm{tr}})=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|$, where $\mathbb{E}_{\widehat{S}_{\mathrm{tr}}}$ is the empirical expectation for $S_{\mathrm{tr}}$. We can derive a similar equation for $S_{\mathrm{te}}$. By using the triangular inequality and setting $l_{i}(Z)=(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}$, we have
$|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)-\mathbb{E}_{Z\sim\widehat{S}_{\mathrm{tr}}}l_{i}(Z)\right|.$ (8)
$\log\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}},\pi}e^{\lambda\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)-\mathbb{E}_{Z\sim\widehat{S}_{\mathrm{tr}}}l_{i}(Z)\right|}\leq B\log 2+4\lambda^{2}/n$ (9)
for any $\lambda\in\mathbb{R}$. Combining this result with the Donsker–Varadhan inequality leads to Eq. (7). ∎
The complete proof is provided in Appendix A.2. This theorem shows that the generalization properties in the ECE behave similarly for the binary and multiclass cases. This is natural, given that the binning-based ECE divides the top-$1$ predicted probabilities equally into $B$ bins in both cases.
Discussion about the assumption of $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$: Although a single data point can be used to evaluate loss in a standard generalization error analysis, multiple data points are required to evaluate the ECE, as it is an estimator. Therefore, for a fair comparison, the same number of data points should be used in the analysis of the ECEs based on $S_{\mathrm{tr}}$ and $S_{\mathrm{te}}$. In this case, it seems difficult to apply the common method of deriving tighter PAC-Bayes bounds, which involves analyzing the KL divergence of the Bernoulli random variable $kl$ [20, 7]. This is because the ECE at $S_{\mathrm{te}}$ is no longer a sum of i.i.d. random variables as it is in a test loss.
Comparison with existing work: The most closely related analysis is provided by Gupta and Ramdas [11]. By comparison, our analysis provides a more general and tighter bound, as explained below. First, they analyzed the ECE for binary classification under the setting where an equal number of samples are allocated in each bin. In contrast, our analysis covers both binary and multiclass classifications under the equal-width binning setting, which is more common in practice. Moreover, their analysis does not include the generalization error analysis, but they rather derived the following concentration bound of the ECE: $|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{val}})|\leq\mathcal{O}_{p}\left(\sqrt{B\log B/n}\right)$, which corresponds to the case where $S_{e}=S_{\mathrm{val}}\sim\mathcal{D}^{n}$ as in our study. Furthermore, when setting $\lambda=\sqrt{BN}$, our bound exhibits $\mathcal{O}_{p}(\sqrt{B/n})$ that improves the order with $\log B$. IN their proof strategy, they evaluate $y-f_{w}(x)$ for each bin, necessitating multiple applications of concentration inequalities. Therefore, aggregating these errors across all bins using a union bound results in a slow convergence. When we adopt this proof technique in our analysis, the resulting bound exhibits $\mathcal{O}_{p}(\sqrt{B\times\mathrm{KL}(\rho\|\pi)/n})$, which is significantly larger than the order of Eq. (7). In comparison, our bound simultaneously treats all bins using Eq. (8) and we should only use Eq. (9) once, resulting in an improved order.
3.2 Estimation bias analysis for calibration
Next, we show the result of our estimation bias analysis under the following smoothness assumption, which is commonly used in the nonparametric estimation [30].
Assumption 1.
Conditioned on $W=w$, the $L$-Lipschitz continuity holds for $\mathbb{E}[Y|f_{w}(x)]$ in the binary classification and $\mathbb{E}[Y|f_{w}(x)_{C}]$ in the multiclass classification.
@@ 3.2 Estimation bias analysis for calibration
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{W},S_{\mathrm{tr}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (10)
Exactly the same statement holds for $\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})$ (see Appendix A.3 for the statement).
Proof sketch.
To analyze the bias caused by binning, following Kumar et al. [18], we define the conditional expectation of $f_{w}$ given bins as $f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}$. With this definition, we decompose the estimation bias as follows: $|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}}^{C})|+|\mathrm{TCE}(f_{\mathcal{I}}^{C})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$. We then evaluate these terms separately. The first term on the right-hand side is the bias caused by binning, which can be bounded using Assumption 1. The second term can be upper-bounded by $\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})$ using Jensen’s inequality. See Appendix A.3 for details. ∎
According to this theorem, the number of bins shows a trade-off relationship as follows. As we increase $B$, the bias caused by binning decreases because we estimate the label frequency more precisely. On the other hand, $\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})$ increases. From this observation, we derive the optimal number of bins that minimizes the estimation bias. Assuming that $\mathrm{KL}(\rho\|\pi)$ is sufficiently smaller than $n$, for example, $\mathcal{O}(\log n)$ and independent of $B$, and setting $\lambda=\sqrt{BN}$, we can minimize the upper bound of Eq. (10) with respect to $B$. This yields $B=\mathcal{O}(n^{1/3})$, and under this bin size, the order of the estimation bias becomes
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{W})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|=\mathcal{O}(\log n/n^{1/3}).$ (11)
Similarly, we can also obtain the same optimal bin size for binary classification.
Relationship with nonparametric regression: The order of the estimation bias in Eq. (11) is tight from the the nonparametric regression viewpoint. Recall that the TCE measures the error between two function: $\mathbb{E}[Y|f_{w}(x)]$ and $f_{w}(x)$. We estimate $\mathbb{E}[Y|f_{w}(x)]$ by binning, which is a nonparametric method. Thus, this is a problem of nonparametric regression on $[0,1]$ under Lipschitz continuity. According to Tsybakov [30], the error in nonparametric regression cannot be smaller than $\mathcal{O}(1/n^{1/3})$. Achieving an order smaller than this necessitates additional assumptions about the underlying data distribution. In this regard, the order of our bound is convincing under the current assumptions. We note that assuming Hölder continuity instead of Assumption 1 does not improve the order of the estimation bias owing to the bias caused by binning (see Appendix A.4 for details).
3.3 Discussion about top-$K$ calibration
In applications such as medical diagnosis [14], it is often important to calibrate the prediction probabilities for any $k\in K$, not simply focusing on the top-$1$. In this case, the alternative metric $CE_{K}\coloneqq\mathbb{E}\|P(\hat{Y}|f_{w}(X))-f_{w}(X)\|_{1}$ has been explored instead of TCE [9], where $\|\cdot\|_{p}$ is the $L^{p}$ distance in $\mathbb{R}^{K}$. The $CE_{K}$ metric measures the calibration performance of models across all classes. Here, we show that the direct estimation of $CE_{K}$ by binning results in a slow convergence due to the curse of dimensionality.
Let us consider the following binning scheme to estimate $CE_{K}$. We split each dimension $[0,1]$ into $B^{\prime}$ bins of the same width because we want to estimate the $K$-dimensional conditional probability in $[0,1]^{K}$ by binning. After doing so, there are $B=(B^{\prime})^{K}$ bins, and $[0,1]^{K}$ is split into $B$ small regions. We refer to these regions as $\mathcal{I}\coloneqq\{I_{i}\}_{i=1}^{B}$. In this case, the ECE of $CE_{K}$ can be expressed as $\mathrm{ECE}_{K}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}\frac{\#B_{i}}{n_{e}}\|\bar{f}_{i,S_{e}}-\bar{y}_{i,S_{e}}\|_{1}$, where #$\#B_{i}$ is the number of instances in $I_{i}$, $\bar{f}_{i,S_{e}}$ is the empirical mean of the prediction, and $\bar{y}_{i,S_{e}}$ is the empirical mean of the one-hot encoding of the label in $I_{i}$. Under these settings, the upper bound of the estimation bias in $\mathrm{ECE}_{K}$ is given as follows.
Theorem 4.
@@ 3.3 Discussion about top-$K$ calibration
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$ and $f_{w}(x)$ has the probability density over $[0,1]^{K}$ and $\mathbb{E}[\hat{Y}|f_{w}(x)]$ satisfies $L$-Lipschitz continuity. Then, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|CE_{K}(f_{W})-\mathrm{ECE}_{K}(f_{W},S_{\mathrm{tr}})|\leq\frac{K(1+L)}{B^{\frac{1}{K}}}+\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{K^{2}\lambda^{2}}{2n}}{\lambda}.$ (12)
@@
By minimizing the upper bound of Eq. (12), we obtain the optimal bin size as $B=\mathcal{O}(n^{\frac{1}{K+2}})$, leading to an estimation bias of $\mathcal{O}(K^{\frac{3}{2}}\log n/n^{\frac{1}{K+2}})$. Although this order aligns with the lower bound of $K$-dimensional nonparametric regression with Lipschitz continuity [30], it is significantly larger than that of the TCE due to the curse of dimensionality. On the other hand, the TCE successfully circumvents this issue as described in Theorem 3 by focusing only on the top-$1$ predicted label, although it does not necessarily guarantee good calibration performance for all $K$ [9].
@@ 4 PAC-Bayes bounds under recalibration
In this section, we extend the bounds from Section 3.1 to the recalibration setting, leading to our novel recalibration algorithm in Section 4.2. All the proofs are provided in Appendix B.
4.1 Bounds for ECE and estimation bias under recalibration
@@
$\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{ECE})\leq\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}$ (13)
This result implies that the KL regularization over the parameter space is required to prevent overfitting similarly to the PAC-Bayes bound when using the training dataset [21, 1]. Since the recalibration data is available and we have fixed $w$, we only consider the posterior $\tilde{\rho}$ over $V$. Instead of fixing $W=w$, we can also obtain the PAC-Bayes bound by taking expectation over $W$ (see Appendix B.2).
Comparison with Theorem 2: $\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})$ is usually smaller than $\mathrm{KL}(\rho\|\pi)$ because of the following two reasons. First, the dimension $d^{\prime}$ of $V$ is much smaller than $d$. This is because, for example, $d^{\prime}$ typically coincides with the number of labels $K$ in the multiclass classification, whereas $d$ is the number of parameters $W$ in large models such as DenseNet [13] used in Section 6. Second, the prior $\tilde{\pi}$ can be dependent on $S_{\mathrm{tr}}$ since $S_{\mathrm{re}}$ and $S_{\mathrm{tr}}$ are independent, which means that we can use a data-dependent prior. These facts suggest that recalibration can improve generalization in terms of the ECE.
@@
Conditioned on $W=w$ and $V=v$, the $L$-Lipschitz continuity holds for $\mathbb{E}[Y|\eta_{v}(f_{w}(X))]$ in binary classification and $\mathbb{E}[Y|\eta_{v}(f_{w}(X))_{C}]$ in the multiclass classification.
@@
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$. Under Assumption 2, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}$ (14)
Exactly the same statement holds for $\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{CE})$ (see Appendix B.1).
Similarly to Eq. (11), there is a trade-off relationship with respect to $B$ in the above bound. By minimizing this upper bound with respect to $B$, we obtain the optimal bin size as $B=\mathcal{O}(n^{1/3})$. For practical purposes, it is possible to set $B=\lfloor n^{1/3}\rfloor$, where $\lfloor x\rfloor\coloneqq\max\{m\in\mathbb{Z}:m\leq x\}$.

### v2

@@ 3.2 Total Bias Under the Training Dataset
The following theorem presents the first PAC-Bayes bound for the generalization error of the ECE.
@@
Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{W},S_{\mathrm{tr}},\mathrm{TCE})$ (11)
$\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n_{\mathrm{tr}}}}{\lambda}.$ (12)
The complete proof is shown in Appendix B.3. Assuming that $\mathrm{KL}(\rho\|\pi)$ is sufficiently smaller than $n_{\mathrm{tr}}$, e.g., $\mathcal{O}(\log n_{\mathrm{tr}})$ and independent of $B$, and setting $\lambda=\sqrt{Bn_{\mathrm{tr}}}$, we can minimize the upper bound of Eq. (11) with respect to $B$. This yields $B=\mathcal{O}(n_{\mathrm{tr}}^{1/3})$ and we have $\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})=\mathcal{O}(\log n_{\mathrm{tr}}/n_{\mathrm{tr}}^{1/3})$.
Comparison with existing work: The most closely related analysis is that performed by Futami and Fujisawa (2024); they evaluated a similar bias using information-theoretic generalization error analysis (IT analysis) (Xu and Raginsky, 2017). Their obtained bound and ours achieved the same order with respect to $n$, which improves previous results, such as those obtained by Gupta and Ramdas (2021). However, our result has two advantages over that obtained by Futami and Fujisawa (2024). The first advantage is that ours can treat the multiclass setting and the second advantage is that our theory builds on the PAC-Bayesian theory and the upper bound is optimizable, which leads to our novel recalibration algorithm, as shown in Section 4. On the other hand, Futami and Fujisawa (2024) used the supersample setting of the IT analysis, with which the optimizable upper-bound is difficult to obtain.
We also note that it seems difficult to use the approach of deriving tighter PAC-Bayes bounds on the basis of the KL divergence of the Bernoulli random variable $kl$ (Maurer, 2004; Foong et al., 2021). This is because the ECE is no longer a sum of i.i.d. random variables as it is in a test loss.
3.3 Analysis for the Top-$K$ Calibration Metric
In applications such as medical diagnosis (Jiang et al., 2012), it is often important to calibrate the prediction probabilities for any $k\in[K]$, not simply focusing on the top-$1$. In this case, the alternative metric
$\mathrm{CE}_{K}(f_{w})\coloneqq\mathbb{E}\|P(Y|f_{w}(X))-f_{w}(X)\|_{1},$ (13)
has been explored instead of TCE (Gruber and Buettner, 2022), where $\|\cdot\|_{p}$ is the $L^{p}$ distance in $\mathbb{R}^{K}$. The $\mathrm{CE}_{K}$ metric measures the calibration performance of models across all classes. Here, we show that the direct estimation of $\mathrm{CE}_{K}$ by binning results in a slow convergence due to the curse of dimensionality.
Let us consider the following binning scheme to estimate $\mathrm{CE}_{K}$. Since we want to estimate the $K$-dimensional conditional probability in $[0,1]^{K}$ by binning, we split each dimension $[0,1]$ into $B^{\prime}$ bins of the same width. After doing so, there are $B=(B^{\prime})^{K}$ bins, and $[0,1]^{K}$ is split into $B$ small regions. We refer to these regions as $\mathcal{I}\coloneqq\{I_{i}\}_{i=1}^{B}$. In this case, the ECE of $\mathrm{CE}_{K}$ can be defined as
$\mathrm{ECE}_{K}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}\|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}\|_{1},$ (14)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}$, $\hat{p}_{i}\coloneqq|I_{i}|/n_{e}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\mathbf{e}_{y_{m}}$. The definition of Eq. (14) is similar to that of Eq. (2). The difference is that Eq. (2) only focuses on the top label. Under these settings, the upper bound of the total bias in $\mathrm{ECE}_{K}$ is given as follows.
@@ 3.3 Analysis for the Top-$K$ Calibration Metric
Assume that $f_{w}(x)$ has a probability density over $[0,1]^{K}$ and $\mathbb{E}[\mathbf{e}_{Y}|f_{w}(x)]$ satisfies the $L$-Lipschitz continuity. Then, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and the posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|\mathrm{CE}_{K}(f_{w})-\mathrm{ECE}_{K}(f_{W},S_{\mathrm{tr}})|$ (15)
$\leq\frac{K(1+L)}{B^{\frac{1}{K}}}+\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{K^{2}\lambda^{2}}{2n_{\mathrm{tr}}}}{\lambda}.$ (16)
By minimizing the upper bound of Eq. (15), we obtain the optimal bin size as $B=\mathcal{O}(n_{\mathrm{tr}}^{\frac{1}{K+2}})$, leading to an estimation bias of $\mathcal{O}(K^{\frac{3}{2}}\log n_{\mathrm{tr}}/n_{\mathrm{tr}}^{\frac{1}{K+2}})$. We can similarly show the result of the total bias as
$\mathrm{Bias}(f_{w},S_{\mathrm{te}},\mathrm{CE}_{K})$ $\coloneqq|\mathrm{CE}_{K}(f_{w})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{te}})|$ (17)
$=\mathcal{O}(1/n_{\mathrm{te}}^{1/(K+2)}).$ (18)
@@
The proofs are provided in Appendix B.8. This order aligns with the lower bound of $K$-dimensional nonparametric regression with Lipschitz continuity (Tsybakov, 2008), and it is much larger than that of the TCE ($\mathcal{O}(\log n_{\mathrm{tr}}/n_{\mathrm{tr}}^{1/3})$) owing to the curse of dimensionality. Thus, this result clarifies the pros and cons of the TCE; the TCE successfully circumvents the curse of dimensionality as described in Theorem 2 by focusing only on the top-$1$ predicted label, although it does not necessarily guarantee good calibration performance for all classes (Gruber and Buettner, 2022).
We note that our proof technique can be extended to cases where we focus on the calibration of any $K^{\prime}\in[1,K]$ classes, resulting in a total bias of $\mathcal{O}(1/n_{\mathrm{te}}^{1/(K^{\prime}+2)})$ (see Appendix B.10 for the details). This finding suggests that in scenarios with a large number of classes, it is crucial to selectively choose which classes to evaluate for calibration, especially when additional classes beyond the top label require calibration evaluation, rather than evaluating all classes. This approach is essential for achieving sample-efficient calibration evaluations.
@@ 4 PAC-Bayes bounds under Recalibration
In this section, we analyze the recalibration using PAC-Bayesian theory, leading to our novel recalibration algorithm in Section 4.2. All the proofs are provided in Appendix C.
4.1 Generalization and Total Bias under Recalibration
@@
$\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{V}\circ f_{w},S_{\mathrm{re}},\mathrm{ECE})$ (19)
$\leq\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (20)
This result highlights the importance of KL regularization in the parameter space—similar to the standard PAC-Bayes bound over $S_{\mathrm{tr}}$ (McAllester, 2003; Alquier et al., 2016)—in preventing overfitting and improving generalization. Since the recalibration data is available and we have fixed $w$, we only consider the posterior $\tilde{\rho}$ over $V$. Instead of fixing $W=w$, we can also obtain the PAC-Bayes bound by taking expectation over $W$ (see Appendix C.2).
Discussion about the assumption of $n_{\mathrm{te}}=n_{\mathrm{re}}=n$: Although a single data point can be used to evaluate loss in a standard generalization error analysis, multiple data points are required to evaluate the ECE, as it is a nonparametric estimator. Therefore, for a fair comparison, the same number of data points should be used in the analysis of the ECEs based on $S_{\mathrm{re}}$ and $S_{\mathrm{te}}$. We also remark that this assumption is not very restrictive since we split all the available data into training, recalibration, and test datasets, and most data are allocated to the training dataset in practice.
@@
Conditioned on $W=w$ and $V=v$, the $L$-Lipschitz continuity holds for $P(Y=C|\eta_{v}\circ f_{w}(X)_{C})$ in the multiclass classification.
@@
Under Assumption 2, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}\circ f_{w},S_{\mathrm{re}},\mathrm{TCE})$ (21)
$\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}.$ (22)
Similarly to Eq. (11), there is a trade-off relationship with respect to $B$ in the above bound. By minimizing this upper bound with respect to $B$, we obtain the optimal bin size as $B=\mathcal{O}(n_{\mathrm{re}}^{1/3})$. For practical purposes, it is possible to set $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$, where $\lfloor x\rfloor\coloneqq\max\{m\in\mathbb{Z}:m\leq x\}$.